OpenStudy (anonymous):

what does for example 6n mean?

OpenStudy (anonymous):

6n means that n is a variable that is being mulitpled by 6

OpenStudy (anonymous):

for example 6(9) is going to be 6 x 9 which is 54

OpenStudy (anonymous):

what would I put when the question is "for any whole number,n,the value of 6n-1 is always a prime number explain how this is wrong?

jhonyy9 (jhonyy9):

so what sign whole number ,,n" -- i think every natural numbers so than if you add value for n in case of 6n-1 you need getting always a prime number so for n=1 result 6*1 -1=5 for n=2 --- 6*2 -1=11 for n=3 --- 6*3 -1=17 for n=4 --- 6*4 -1=23 for n=5 --- 6*5 -1=29 for n=6 --- 6*6 -1 =35 so what not is a prime number si than result that this statement is wrong